# Distance-coupling as an Approach to Position and Formation Control

### v2

@@ III-D Distance-coupled Homing Results
Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately for $x^{(0)}$ chosen outside $Rx^{(\text{eq})}+r$.
Figure 4: Ratio of $100$ tests diverged for $100$ rotations, defined by $R:\theta\in[-2\pi,2\pi]$, of the anchor positions in $\mathcal{A}$.
We now validate the response of the distance-coupled controller (13) with varying degrees of noise in the measurement terms. The claims made on the regions of attraction defined in (22) will also be demonstrated. Note that the control coefficients are defined as $C=\alpha\mathbf{I}:\alpha=-5$ in all cases.
Let us define $\omega_{\varepsilon}$ as a uniformly distributed random variable with expected values in the range $[-\varepsilon,\varepsilon]$. The measurement noise is thus described as
$d_{i}(x,\varepsilon)=d_{i}(x)+\omega_{\varepsilon}.$
Figure 2 shows the response of the system with $p=3$ anchors and for varying degrees of noise, $\varepsilon$. The first case shows the ideal measurement scenario. That is, $\varepsilon=0$ with no error in the positions of the anchors. We place $x^{(\text{eq})}$ at the origin for simplicity. In this case, it is clear that there is no deviation between the true position and the anchor frame approximation. Likewise, the Lyapunov function converges ($V(x)\rightarrow 0$) as predicted from Proposition 1.
With an incorporation of a relatively large degree of measurement error ($\varepsilon=5$ and $\varepsilon=10$) we observe considerable differences between the approximation of the position from the distance-coupled function and the true position of the agent. Nonetheless, the agent is still capable of converging to a region around the equilibrium from any point.
Next, we validate the claims made in Section III-C for transformations of the anchors without updating (11). We first define the rotation $R:\theta=\tfrac{\pi}{4}$ and show that the agents converge with a policy defined by $R^{\intercal}C$. In the translation case, we sample various $r$ as uniform translations of the anchor set. Both sets of results are shown in Figure 4.
To demonstrate the region of attraction with respect to the rotation, $R$, of the anchors we can look at the behavior of the system for various angles $\theta$. Figure 4 shows the divergence rate for an agent with initial conditions
$x^{(0)}=x^{(\text{eq})}+\omega_{\varepsilon},$
where $\varepsilon=0.001$ is intentionally selected to be very small. We then show the results for $100$ uniformly spaced $\theta\in[-2\pi,2\pi]$ which were simulated until the candidate function converged ($V(x)\rightarrow 0$) or diverged ($V(x)\rightarrow\infty$). Each value in Figure 4 is the result of $100$ initial conditions. The system clearly stabilizes for all $\theta$ defined by the bounds in (22), and diverges otherwise. For a more explicit example, Figure 5 shows a series of agents initialized with $\omega_{\varepsilon=2.5}$ and $R:\theta=\tfrac{3\pi}{4}$ diverging from the origin as expected from (22).
@@ IV Distance-coupled Formation Control
Using the distance-coupled homing controller (13), we will define a new policy which drives $m$-agents to some predetermined formation. The only measurements each agent can make is the Euclidean distance it stands from each of the other agents in the environment. The following sections discuss the new system’s structure, defines the distance-coupled formation controller and proves certain properties analytically before demonstrating said claims in simulation.
IV-A Distance-coupled Formation Control
We utilize $m=p$ agents, with positions denoted by (3) and an intended formation described by the set $\mathcal{A}$. By defining our formation in terms of $\mathcal{A}$ we can exploit the notation from (12) to approximate the position of the agents in the anchor frame (discussed in Section IV-B).
In this scenario, each agent acts as an anchor for the others and vice versa. Thus, for an agent, $x_{k}$, we have the agent-distance set, $\mathcal{D}(x_{k},X)$, which is defined similarly to (6). We note that, in this form, the $k$-th element of the agent-distance set is always equal to 00.
@@ IV-A Distance-coupled Formation Control
where hstack()($\cdot$)( ⋅ ) is the horizontal alternative to vstack()($\cdot$)( ⋅ ). In order to utilize (12) for all agents, we expand the measurement vector from (11) to the matrix:
$H(X)=\begin{bmatrix}h(x_{1})&\cdots&h(x_{m})\end{bmatrix},$ (27)
where $h(x_{k})$ is the distance-coupled agent measurements derived from $\mathcal{D}(x_{k},X)$, i.e.
@@
We similarly arrange the squared anchor difference vector from (11) as
$B=b\mathbf{1}_{m}^{\intercal},$ (29)
where $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of all ones. This is valid as the vector $b$ is equivalent for every agent in $X$. Using the matrix notations defined above we can reformulate (12) and (13) as
@@
where $K$ is similarly unchanged from (12). Finally, we can note that the only parameter which differs between the agents is the intended equilibrium position, identified by $x^{(\text{eq})}_{k}$ in the matrix $X^{(\text{eq})}$.
@@ IV-B Formation Equivalence in the Anchor Frame
Before further inspection, we define formation equivalence (or, congruence) between the two frames of reference, $\mathbb{A}$ and $\mathbb{W}$, using selected tools from Procustes analysis [15]. Similar to the offsets defined in Section III-C, we have the agent positions in the world frame, $X\in\mathbb{W}$, with corresponding anchor frame coordinates $\tilde{X}\in\mathbb{A}$.
@@
Procrustes analysis defines the two sets, $X$ and $\tilde{X}$, as congruent if
@@
given the orthogonal rotation, $\Psi(X,\tilde{X})=R(\theta)$, and the centroid of the formation, $\langle x\rangle=\tfrac{1}{m}\Sigma_{k=1}^{m}x_{k}$.
In practice, we can calculate an approximation of $\Psi$ using the Kabsch algorithm [16]. The resulting matrix is the most optimal rotation which superposes the set $\tilde{X}$ onto $X$.
Another method of denoting the transformation (32) is to combine the constant terms on the left and right sides into a single linear translation which superposes the anchor frame formation onto the world frame. This turns (32) into
$X-\langle x\rangle=\Psi(\tilde{X}-\langle\tilde{x}\rangle)\;\Rightarrow\;X=\Psi\tilde{X}+\psi,$ (33)
where $\psi=\langle x\rangle-\Psi\langle\tilde{x}\rangle$. The later sections will primarily use the shorthand definition of formation error (33). It can be further noted that the uniform scaling component to Procrustes analysis is omitted.
In order to measure policy convergence in Section IV-E, we define the formation error between $X$ and $X^{(\text{eq})}$ as the $L_{2}$-norm of the function (33) on individual agents, i.e.
@@ IV-C Transformation Invariance
A function $f(x,y):x,y\in\mathbb{R}^{n}$ is transformation invariant if $f(x,y)$ is constant under equivalent rotations and translations of both $x$ and $y$. That is,
@@
given that $R\in\mathbb{R}^{n\times n}$ is an orthogonal rotation and $r\in\mathbb{R}^{n}$.
@@
Define an arbitrary orthogonal rotation, $R$, and translation, $r$, of the original states $x$ and $y$ such that
$x\rightarrow Rx+r\text{, and }y\rightarrow Ry+r.$ (37)
Taking the distance between the transformed points:
$d(Rx+r,Ry+r)=\\ \sqrt{(Rx+r-(Ry+r))^{\intercal}(Rx+r-(Ry+r))}\\ \begin{aligned} &=\sqrt{(x-y)^{\intercal}R^{\intercal}R(x-y)}\\ &=\sqrt{(x-y)^{\intercal}(x-y)},\end{aligned}$
which is equivalent to (36). Thus concluding that the distance function follows by Definition 2.
@@ IV-D Equilibria of the Distance-coupled Formation Controller
Figure 6: Distance-coupled formation control response for an arbitrary formation pattern and with varying levels of randomly generated offsets from the equilibrium positions: $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon}$.
Using Lemma 2 we establish the invariance of the distance-coupled functions in $H(X)$. Subsequently, we will use this claim to derive the equilibria of (31).
@@
The distance-coupled position (30) is transformation invariant for any congruent set of positions $X=\Psi\tilde{X}+\psi$.
@@
Let there exists an orthogonal rotation $\Psi$ and translation $\psi$ so that $X$ and $\tilde{X}$ are congruent. Identifying that the dependent variables in (30) are solely contained in $H(X)$, we isolate a single row of the measurement terms for an arbitrary agent, $h(x_{k})$. That is,
@@
for given indices $(i,j)$. Applying the transformation $(\Psi,\psi)$ to the entire formation, we get
$h_{i,j}(\Psi\tilde{x}_{k}+\psi)=d^{2}(\Psi$ $\tilde{x}_{k}+\psi,\Psi\tilde{x}_{i}+\psi)$ (39)
$-d^{2}(\Psi\tilde{x}_{k}+\psi,\Psi\tilde{x}_{j}+\psi).$
Identifying the transformation invariance of the distance function from Lemma 2:
$h_{i,j}(\Psi\tilde{x}_{k}+\psi)$ $=d^{2}(\tilde{x}_{k},\tilde{x}_{i})-d^{2}(\tilde{x}_{k},\tilde{x}_{j})$ (40)
$=h_{i,j}(\tilde{x}_{k}).$
@@
Proving that (30) is transformation invariant.
@@
The controller (31) is transformation invariant for any congruent $X$ and $\tilde{X}$.
@@
Let us assume that Proposition 3 is true for the rotation $\Psi$ and translation $\psi$. Then we have that
@@
$=C(K(H(\tilde{X})-B)-X^{(\text{eq})})$
$=U(\tilde{X}).$
@@
We can then use Proposition 4 to identify the equilibria of the distance-coupled formation controller.
@@
Proof.
Assume that $\tilde{X}=X^{(\text{eq})}$ such that $U(\tilde{X})=0$. Then, for the world frame positions, $X=\Psi\tilde{X}+\psi$, and by Proposition 4 we have that
$U(X)=U(\Psi\tilde{X}+\psi)=U(\tilde{X})=U(X^{(\text{eq})})=0.$ (43)
Proving that the positions $X$ are at an equilibrium given they are congruent to $\tilde{X}$.
It should be noted that, while Corollary 1 defines the equilibria of (31), it makes no claims on its stability. However, because of the similarities between (13) and (31), it can be reasonably deduced that, for agent positions $X$ which are not congruent to $\tilde{X}$, we get similar bounds on the region of attraction as those shown in (22). This is demonstrated in the results for small perturbations of the agents from the equilibrium for various rotations.
IV-E Distance-coupled Formation Results
Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
Figure 7: Response of formation controller with offset $\omega_{\varepsilon=10}$ and for the two agent sets with positions denoted by $X_{2}=X_{1}+r$.
Figure 8: Region of attraction results with $p=4$ and $100$ repetitions at each $R:\theta\in[-2\pi,2\pi]$ for a randomly generated offset with $\omega_{\varepsilon=0.1}$.
While we do not present a proof for the stability of the distance-coupled formation controller, we can demonstrate empirically its behavior under various initial conditions. We first show the response of the policy when the initial conditions are randomly offset around $X^{(\text{eq})}$. This will also help to validate our claims on the equilibrium positions defined in Corollary 1. The region of attraction will then be plotted similarly to Figure 4, and the results will be compared.
To validate the response of (31) we initialize the agent group such that
$X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon},$
with $\varepsilon=10$ and $\varepsilon=25$. The simulation is then iterated and the behavior of the policy is shown. The performance of the formation error function $W(X)$ is plotted over time for each test, showing that it converges in these two cases.
Figure 8 uses two sets of agents to validate the convergence of the formation controller for similarly perturbed formations. The first set, $X_{1}$, is the reference set and has initial conditions which are generated similar to the Figure 6 example with $\varepsilon=10$. The second set is linearly offset from $X_{1}$ by a constant $r$. As expected, the two sets follow the same trajectory and end in equilibrium positions which differ by the initial translation, $r$.
Because of their close relationship, we tested if the controller would behave similarly to rotations about the origin as it does in the homing problem. We thus modified the initial condition formula for the agent set such that
$X^{(0)}=R(X^{(\text{eq})}+\omega_{\varepsilon}),$
with $\varepsilon=0.1$ and for the same list of $\theta$ described in Section III-D. The simulation was then iterated until the controller converged ($W(X)\rightarrow 0$) or diverged ($W(X)\rightarrow\infty$), and the proportion of tests which diverged for $100$ samples are shown in Figure 8. As expected, there is a strong correlation to the bounds defined in (22), with more complicated interplay in the formation policy case. Figure 9 shows an example of the controller diverging with $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
V Conclusion
Figure 9: Divergence of (31) with initial conditions $x^{(0)}=Rx^{(\text{eq})}$ given $R:\theta=\pi$ and $\omega_{\varepsilon=0.1}$.
In this paper we define the distance-coupling method as a suitable approach to the position and formation control problems. In the homing case we prove stability and show that it is held for linear transformations of the anchor positions, even when the policy components (11) are not updated. The distance-coupled formation controller is also studied, and the equilibrium conditions are established under the definitions of congruence and transformation invariance. Claims are made on the general performance of the system under arbitrary transformations, and the resulting convergence properties are demonstrated empirically.
Generally speaking, the approach described here is framed using a single-integrator model (1) for simplicity. That said, the distance-coupled position functions (12) and (30) can be applied to any localization problem in which the available measurements are a function of the distance from known reference points. In the context of systems with more complex dynamics, the distance-coupled policies defined here can be used to calculate suitable waypoints for the agents to follow by a more appropriate controller.

### v3

@@ III-D Distance-coupled Homing Results
We now validate the response of the distance-coupled controller (8) with varying degrees of noise in the measurement terms and different angles $\theta$ between $\mathbb{A}$ and $\mathbb{W}$. The claims made on the regions of attraction defined in (10) will also be demonstrated. Note that the control coefficients are defined as $C=\alpha\mathbf{I}$ where $\alpha=-5$.
We define $\omega_{\varepsilon}\in[-\varepsilon,\varepsilon]$ as a uniformly distributed random variable. The measurement noise is thus described as $d_{i}(x,\varepsilon)=d_{i}(x)+\omega_{\varepsilon}$.
Figure 2 shows the response of the system with $p=3$ anchors and varying $\varepsilon$. In all tests we place $x^{(\text{eq})}$ at the origin. The first case shows the ideal measurement scenario, $\varepsilon=0$. It is clear that there is no deviation between the true position and the anchor frame approximation. Likewise, the Lyapunov function converges ($V(x)\rightarrow 0$) as predicted from Proposition 1. With a relatively large degree of measurement error ($\varepsilon=1$ and $\varepsilon=2.5$) we observe considerable differences between the anchor frame approximation and the true position. That said, the agent still converges to a region around the equilibrium from any point.
Next, we validate the claims made in Section III-C for transformations of the anchors without updating the policy components (6). We first define the rotation $R$ with $\theta=\tfrac{\pi}{4}$, and show that the agents converge with a policy defined by $R^{\intercal}C$. We then demonstrate the effects of various translations $r$ over the anchor set and show that $x\rightarrow x^{(\text{eq})}+r$ as the simulation progresses. Both cases are shown in Figure 3. To demonstrate the instability defined in Proposition 2 we rotate the anchors by $\theta=\tfrac{3\pi}{4}$ which does not fall in the range defined in (10) with $k=0$ (Figure 4).
Figure 3: Homing controller response with $x^{(\text{eq})}=0$ and with variations in $R$ and $r$ simulated separately with (left) initial conditions around the origin, and (right) $x^{(0)}=x^{(\text{eq})}$.
@@ IV Distance-coupled Formation Control
Using the distance-coupled homing controller (8), we will define a new policy which drives $m$-agents to some predetermined formation. The only measurements each agent can make is the Euclidean distance it stands from each of the other agents in the environment. The following sections discuss the new system’s structure, define the distance-coupled formation controller, and prove certain properties analytically before demonstrating in simulation.
IV-A Distance-coupled Formation Functions
We utilize $m=p$ agents, with positions denoted by $X$ and an intended formation described by the set $\mathcal{A}$. By defining our formation in terms of $\mathcal{A}$ we can use the notation from (7); approximating the positions of the agents in the anchor frame as discussed in Section IV-B.
In this scenario, each agent acts as an anchor for the others and vice versa. Thus, for an agent $x_{k}$, we have the agent-distance set $\mathcal{D}(x_{k},X)$, defined similarly to (2). We note that the $k$-th element of $\mathcal{D}(x_{k},X)$ is always equal to 00 because of its inclusion in $X$.
@@ IV-A Distance-coupled Formation Functions
where hstack()($\cdot$)( ⋅ ) is the horizontal concatenation function. In order to use (7) for all agents, we expand the measurement vector from (6) to the matrix $H(X)=[h(x_{1})\cdots h(x_{m})]$, where $h(x_{k})$ is the distance-coupled agent measurements derived from $\mathcal{D}(x_{k},X)$, i.e.
@@
We similarly arrange the squared anchor difference vector from (6) as $B=b\mathbf{1}_{m}^{\intercal}$ where $\mathbf{1}_{m}\in\mathbb{R}^{m}$ is a vector of all ones. This is valid as the vector $b$ is equivalent for every agent in $X$. We can then reformulate (7) and (8) as
@@
where $K$ is unchanged from (7). Finally, we can note that the only parameter which differs between the agents is the intended equilibrium position.
@@ IV-B Formation Equivalence in the Anchor Frame
Before further analysis, we define formation equivalence between the two frames of reference, $\mathbb{A}$ and $\mathbb{W}$, using tools from Procustes analysis. Similar to the offsets defined in Section III-C, we have the agent positions in the world frame, $X\in\mathbb{W}$, with corresponding anchor frame coordinates $\tilde{X}\in\mathbb{A}$.
@@
Procrustes analysis defines the two sets, $X$ and $\tilde{X}$, as congruent if there exists a rotation $\Psi$ such that
@@
where $\langle x\rangle=\tfrac{1}{m}\Sigma_{k=1}^{m}x_{k}$ is the centroid of the formation.
In practice, we can compute an approximation of $\Psi$ using the Kabsch algorithm [kabsch_discussion_1978]. The resulting matrix is the optimal rotation to superpose the set $\tilde{X}$ onto $X$.
The transformation (17) can also be written as $\Psi X+\psi$ where $\psi=\langle x\rangle-\Psi\langle\tilde{x}\rangle$. In order to measure convergence, we define the $L_{2}$-norm of the formation error as
@@ IV-C Transformation Invariance
A function $f(x,y):x,y\in\mathbb{R}^{n}$ is transformation invariant if $f(x,y)$ is constant under equivalent rotations and translations of both $x$ and $y$, that is,
@@
for any orthogonal rotation $R\in\mathbb{R}^{n\times n}$ and offset $r\in\mathbb{R}^{n}$.
@@
Follows trivially from expanding and simplifying $d(Rx+r,Ry+r)$.
@@ IV-D Equilibria of the Distance-coupled Formation Controller
Using Lemma 2 we establish the invariance of the distance-coupled functions in $H(X)$ before deriving the equilibria of (16).
@@
The distance-coupled position (15) is transformation invariant when $X=\Psi\tilde{X}+\psi$ are congruent.
@@
Identifying that the dependent variables in (15) are solely contained in $H(x)$, we isolate a single row of the measurement terms for the $k$-th agent, $h(x_{k})$. This yields,
@@
for given indices $(i,j)$. Applying the transformation $x_{k}=\Psi\tilde{x}_{k}+\psi$ to each agent, and using Lemma 2, we get,
$h_{i,j}(x_{k})$ $=h_{i,j}(\Psi\tilde{x}_{k}+\psi)$ (22)
$=d^{2}(\tilde{x}_{k},\tilde{x}_{i})-d^{2}(\tilde{x}_{k},\tilde{x}_{j})=h_{i,j}(\tilde{x}_{k}).$
@@
The claim then follows for (7).
@@
The controller (16) is transformation invariant when $X$ and $\tilde{X}$ are congruent.
@@
By the same logic as Proposition 3, with rotation $\Psi$ and translation $\psi$, we have that
@@
$=C(K(H(\tilde{X})-B)-X^{(\text{eq})})=U(\tilde{X}),$
@@
We then use Proposition 4 to identify the equilibria of the distance-coupled formation controller.
@@
The formation controller (16) is at an equilibrium, $U(X)=0$, when $X$ is congruent to $X^{(\text{eq})}$.
Proof.
Define $\tilde{X}$ at $X^{(\text{eq})}$ such that $U(\tilde{X})=0$. Then, by Proposition 4 we have that $U(X)=U(\Psi\tilde{X}+\psi)=U(\tilde{X})=0$. Thus, the formation $X=\Psi\tilde{X}+\psi$ is at an equilibrium given it is congruent to $X^{(\text{eq})}$.
It should be noted that, while Corollary 1 defines the equilibria of (16), it makes no claims on its stability. However, because of the similarities between (8) and (16), it can be reasonably concluded that, for agent positions $X$ which are not congruent to $\tilde{X}$, we get similar bounds on the region of attraction as those defined in Lemma 1. This claim is validated empirically in Section IV-E.
IV-E Distance-coupled Formation Results
Figure 5: Distance-coupled formation control response for an arbitrary formation and with varying levels of randomly generated offsets from the equilibrium positions: $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon=25}$.
Figure 6: Response of formation controller with $X_{1}=X^{(\text{eq})}+\omega_{\varepsilon=10}$ and $X_{2}=X_{1}+r$.
Figure 7: Region of attraction results with $p=4$ and $100$ repetitions at each $R$ with $\theta\in[-2\pi,2\pi]$ and for an initial position $X^{(0)}=X^{(\text{eq})}+\omega_{\varepsilon=0.1}$.
While we do not present a proof for the stability of the distance-coupled formation controller, we can demonstrate empirically its behavior under various initial conditions. We first show the response of the policy when the initial conditions are randomly offset around $X^{(\text{eq})}$. This will also help to validate our claims on the equilibrium positions defined in Corollary 1.
To validate the response of (16) we initialize the agent group such that
$X^{(0)}=R(X^{(\text{eq})}+\omega_{\varepsilon}),$
with noise value $\varepsilon=25$, and rotation $R=\mathbf{I}$. The simulation is then iterated and the behavior of the policy is shown. The performance of the formation error function $W(X)$ is plotted over time; showing that it converges in this case.
Figure 6 uses two sets of agents to validate the convergence of the formation controller for similarly perturbed formations. The first set, $X_{1}$, is the reference set and has initial conditions which are generated similar to the Figure 5 example with $\varepsilon=10$. The second set $X_{2}$ is linearly offset from $X_{1}$ by a constant $r$. As expected, the two sets follow the same trajectory and end in equilibrium positions which differ by the initial translation, $r$.
Because of their close relationship, we tested if the controller would behave similarly to rotations about the origin as it does in the homing problem. We thus take $\varepsilon=0.1$ and select $R$ for a uniformly distributed set $\theta\in[-2\pi,2\pi]$. The simulation was then iterated until the controller converged ($W(X)\rightarrow 0$) or diverged ($W(X)\rightarrow\infty$), and the proportion of tests which diverged for $100$ samples are shown in Figure 7. As conjectured, there is a strong correlation to the bounds defined in (10).
V Conclusion
In this paper we define the distance-coupling method as a suitable approach to the position and formation control problems. In the homing case, we prove stability and show that it is held for linear transformations of the anchor positions, even when the policy components (6) are not updated. The distance-coupled formation controller is also studied, and the equilibrium conditions are established under the definitions of congruence and transformation invariance. Claims are made on the general performance of the system under arbitrary transformations, and the resulting convergence properties are demonstrated empirically. A key feature in our approach is that the controller can be implemented without relative position or bearings.
Generally speaking, the approach described here is framed using a single integrator model for simplicity. That said, the distance-coupled position functions (7) and (15) can be applied to any localization problem in which the available measurements are a function of the distance from known reference points. In the context of systems with more complex dynamics, the distance-coupled policies defined here can be used to calculate suitable waypoints for the agents to follow by a more appropriate controller.